# Matryoshka: Learning to Drive Black-Box LLMs with LLMs

### v1

@@ G.1 ALFWorld
Following Adaplanner (Sun et al., 2024a), we implement a code-style prompt for Matryoshka, which can be divided into the following sections:
@@
The <high_level_planning> prompt is used to instruct Matryoshka to break the current task down into multiple subtasks, where <decompose> is replaced by a standard task decomposition process, and <receptacle_list> is substituted by the list of interactive receptacles provided by the task environment. Finally, <task> is replaced by the task description, expressed in natural language.
 <high_level_planning> Prompt
 # Decompose the task into steps. First give a general plan of how you
 would solve the task, then for each step you plan to take, mark with
 ’[Step xx]’.
 # Here is an example of a decomposition to the task:
 # define environment
 receptacles = [’diningtable 1’,’drawer 2’, ’drawer 1’, ’sinkbasin 1’,
 ’toilet 1’, ’sidetable 2’, ’sidetable 1’, ’cabinet 1’, ’countertop 1’,
 ’microwave 1’, ’fridge 1’]
 <decompose>
 # Here is the actual task.
 # define environment
 receptacles = <receptacle_list>
 # <task>
 # here is a decomposition:
@@
The <low_level_execution> prompt is used to instruct the black box model to generate a specific solution based on the problem and the plan provided by Matryoshka. <basic_info> defines the agent and admissible actions on Alfworld and can be found in Sun et al. (2024a). The <example> is replaced with a combination of planning and expert trajectory, while the meanings of <receptacle_list> and <task> remain consistent with the previous description. <decomposition> represents the high-level plan provided by Matryoshka for the current task.
 <low_level_execution> Prompt
 <basic_info>
 # Now complete the function solution() below to solve the task by
 composing the agent’s methods to interact with the environment.
 # First give a general plan of how you would solve the task, mark with
 ’ # General Plan’. Then for each step you plan to take, 1) mark with
 ’[Step xx]’, 2) give a reason why you think it is a good step to take
 3) write an assertion to check if the step is successful.
 # Here is an example of a solution to the task:
 # define environment and agent
 receptacles = [’diningtable 1’,’drawer 2’, ’drawer 1’, ’sinkbasin 1’,
 ’toilet 1’, ’sidetable 2’, ’sidetable 1’, ’cabinet 1’, ’countertop 1’,
 ’microwave 1’, ’fridge 1’]
 agent = Agent(receptacles)
 <example>
 # Here is the actual task.
 # define environment and agent
 receptacles = <receptacle_list>
 agent = Agent(receptacles)
 # <task>
 # here is a decomposition:
 <decomposition>
 # here is a solution:
@@
In ALFWorld, there are six types of tasks: Pick, Clean, Heat, Cool, Examine, and Pick two. For each type, we collect a reasonable high-level planning approach, allowing Matryoshka to reference them. These six planning samples are presented as follows:
@@
 <planning_sample_pick> Prompt
 # Your task is to: put soapbar on countertop.
 # here is a decomposition:
 # General Plan: I need to get a list of receptacles where the soapbar
 is likely to appear, and then go to each receptacle in the list until
 seeing a soapbar. Then I can put get the identifier of the soapbar and
 take it. Finally I can go to the countertop and put the soapbar.
 # [Step 1] get a list of receptacles where the soapbar is likely to
 appear.
 # [Step 2] go to each receptacle in the list until seeing a soapbar.
 # [Step 3] identify the soapbar I juts found and take it.
 # [Step 4] go to a countertop and put the soapbar on it.
@@
 <planning_sample_clean> Prompt
 # Your task is to: put a clean lettuce in diningtable / clean a
 lettuce and put it in diningtable.
 # here is a decomposition:
 # General plan: I need to get a list of receptacles to find the
 lettuce, take the lettuce to the sinkbasin, clean it and put it in a
 diningtable.
 # [Step 1] get a list of receptacles where the lettuce is likely to
 appear.
 # [Step 2] go to each receptacle in the list until seeing a lettuce.
 # [Step 3] identify the lettuce I just found and take it.
 # [Step 4] go to a sinkbasin to clean the lettuce.
 # [Step 5] go to a diningtable and put the lettuce on it.
@@
 <planning_sample_heat> Prompt
 # Your task is to: put a hot lettuce in diningtable / heat some
 lettuce and put it in diningtable.
 # here is a decomposition:
 # General plan: I need to get a list of receptacles to find the
 lettuce, take the lettuce to the microwave, heat it and put it in a
 diningtable.
 # [Step 1] get a list of receptacles where the lettuce is likely to
 appear.
 # [Step 2] go to each receptacle in the list until seeing a lettuce.
 # [Step 3] identify the lettuce I juts found and take it.
 # [Step 4] go to a microwave to heat the lettuce.
 # [Step 5] go to a diningtable and put the lettuce on it.
@@
 <planning_sample_cool> Prompt
 # Your task is to: put a cold lettuce in diningtable / cool some
 lettuce and put it in diningtable.
 # here is a decomposition:
 # General plan: I need to get a list of receptacles to find the
 lettuce, take the lettuce to the fridge, cool it and put it in a
 diningtable.
 # [Step 1] get a list of receptacles where the lettuce is likely to
 appear.
 # [Step 2] go to each receptacle in the list until seeing a lettuce.
 # [Step 3] identify the lettuce I juts found and take it.
 # [Step 4] go to a fridge to cool the lettuce.
 # [Step 5] go to a diningtable and put the lettuce on it.
@@
 <planning_sample_examine> Prompt
 # Your task is to: look at the bowl under the desklamp / examine the
 bowl with the desklamp
 # here is a decomposition:
 # General plan: I need to get a list of receptacles to find the bowl
 and take the bowl with me, then I get another list of receptacles to
 find the desklamp and turn it on.
 # [Step 1] get a list of receptacles where a bowl is likely to appear.
 # [Step 2] go to each receptacle in the list until seeing a pen.
 # [Step 3] take the bowl from the receptacle.
 # [Step 4] get a list of receptacles where a desklamp is likely to
 appear.
 # [Step 5] go to each receptacle in the list until seeing a desklamp.
 # [Step 6] turn on desklamp.
@@
 <planning_sample_picktwo> Prompt
 # Your task is to: put two cellphone in cabinet / find two cellphone
 and put them in cabinet
 # here is a decomposition:
 # General plan: I need to get a list of receptacles to find the two
 cellphones, find and take the first cellphone and put it in a cabinet,
 then find and take the second cellphone and put it in the cabinet.
 # [Step 1] get a list of receptacles where a cellphone is likely to
 appear.
 # [Step 2] go to each receptacle in the list until seeing a cellphone.
 # [Step 3] identify the first cellphone found and take it.
 # [Step 4] go to a cabinet and put the first cellphone found on it.
 # [Step 5] go to each of the remaining receptacle in the list until
 seeing a second cellphone.
 # [Step 6] identify the second cellphone I just found and take it.
 # [Step 7] go to a cabinet and put the second cellphone found on it.
@@
Our execution sample is based on the prompt structure from Sun et al. (2024a), with the key distinction being the incorporation of the planning component. In this setup, <decompose> is substituted with the task-specific planning sample, <execution> is replaced by the expert samples from Sun et al. (2024a), and the definition of <task> remains unchanged from the previous description.
 <execution_sample_template> Prompt
 # <task>
 # <decompose>
 # <execution>
@@
To implement close-loop refinement during the inference stage, we follow the approach from Sun et al. (2024a) and introduce several prompts: a <code_check> prompt to identify and fix any syntax errors during execution generation, a <refinement> prompt to address refinement in case of assertion errors, and a <start_from> prompt to determine the starting point for the new solution after revising the plan. Detailed descriptions of these prompts can be found in Sun et al. (2024a).
G.2 GSM-Hard
Following PAL framework (Gao et al., 2023), we implement a code-based framework to solve mathematical problems on GSM-Hard, which is primarily divided into two steps: Matryoshka breaks down the mathematical problem into sub-problems, and the black-box model converts each sub-problem into a code block.

### v2

@@ G.1 Alfworld
We compare M-Pilot with both the LLaMA3-8B-Instruct model used as the controller and AdaPlanner [41] in terms of performance across various tasks and environments. In M-Pilot, the controller model is trained based on LLaMA3-8B-Instruct, while the black-box executor employed is gpt-3.5-turbo-0125. Our results demonstrate that M-Pilot enhances the black-box executor’s ability to interact with the environment, enabling it to successfully accomplish target tasks. In contrast, both AdaPlanner and the LLaMA3-8B-Instruct model commonly encounter issues such as initiating actions that the agent cannot complete and misjudging the agent’s state. M-Pilot effectively mitigates these problems, resulting in more reliable task execution.
Task Definition You are in the middle of a room. Looking quickly around you, you see a bed 1, a desk 2, a desk 1, a drawer 6, a drawer 5, a drawer 4, a drawer 3, a drawer 2, a drawer 1, a garbagecan 1, a laundryhamper 1, a safe 1, a shelf 6, a shelf 5, a shelf 4, a shelf 3, a shelf 2, and a shelf 1. Your task is to: examine the book with the desklamp.
M-Pilot # Decomposition: General plan: I need to get a list of receptacles to find the book and take the book with me, then I get another list of receptacles to find the desklamp and turn it on. [Step 1] get a list of receptacles where a book is likely to appear. [Step 2] go to each receptacle in the list until seeing a book. [Step 3] take the book from the receptacle. [Step 4] get a list of receptacles where a desklamp is likely to appear. [Step 5] go to each receptacle in the list until seeing a desklamp. [Step 6] turn on the desklamp. [Step 7] go to the receptacle where the book is and turn on the desklamp. [Step 8] examine the book with the desklamp. # GPT Solution: def solution(agent, start_from=1): # General plan: I need to get a list of receptacles to find the book and take the book with me, then I get another list of receptacles to find the desklamp and turn it on. if start_from <= 1: print(”[Step 1] get a list of receptacles where a book is likely to appear.”) # Ask the assistant to sort the receptacles based on the likelihood of finding a book. answer = ask(fGiven a list of receptacles, please sort them in descending order based on the likelihood of finding a book in each of them. The list of receptacles is: {agent.receptacles}. You should directly return a Python list.’) recep_to_check = literal_eval(answer) # Check if the returned list is not empty. assert recep_to_check, fError in [Step 1]: recep_to_check should not be empty. {agent.report()}’ if start_from <= 2: print(”[Step 2] go to each receptacle in the list until seeing a book.”) for receptacle in recep_to_check: observation = agent.goto(receptacle) # Check if the receptacle is closed. If so, open it. if closed in observation: observation = agent.open_receptacle(receptacle) # Check if a book is in/on the receptacle. if book in observation: break # Check if a book is found in one of the receptacles. assert book in observation, fError in [Step 2]: There is no book in/on {recep_to_check}. {agent.report()}’ if start_from <= 3: print(”[Step 3] take the book from the receptacle.”) # Get the identifier of the book. answer = ask(fFrom the observation, get the identifier of an object. For example, On the cabinet 1, you see a cloth 2, and a book 1. The identifier of cloth is 2. Now, {observation} The identifier of the book? Only Output a single number without any other words.’) found_book = fbook {answer}’ observation = agent.take(found_book, receptacle) # Check if the agent is holding the book. assert agent.holding == found_book, fError in [Step 3]: I cannot take {found_book} from the {receptacle}. {agent.report()}’ if start_from <= 4: print(”[Step 4] get a list of receptacles where a desklamp is likely to appear.”) # Ask the assistant to sort the receptacles based on the likelihood of finding a desklamp. answer = ask(fGiven a list of receptacles, please sort them in descending order based on the likelihood of finding a desklamp in each of them. The list of receptacles is: {agent.receptacles}. You should directly return a Python list.’) recep_to_check = literal_eval(answer) # Check if the returned list is not empty. assert recep_to_check, fError in [Step 4]: recep_to_check should not be empty. {agent.report()}’ if start_from <= 5: print(”[Step 5] go to each receptacle in the list until seeing a desklamp.”) for receptacle in recep_to_check: observation = agent.goto(receptacle) # Check if the receptacle is closed. If so, open it. if closed in observation: observation = agent.open_receptacle(receptacle) # Check if a desklamp is in/on the receptacle. if desklamp in observation: break # Check if a desklamp is found in one of the receptacles. assert desklamp in observation, fError in [Step 5]: There is no desklamp in/on {recep_to_check}. {agent.report()}’ if start_from <= 6: print(”[Step 6] turn on the desklamp.”) # Get the identifier of the desklamp. answer = ask(fFrom the observation, get the identifier of an object. For example, On the cabinet 1, you see a cloth 2, and a desklamp 1. The identifier of cloth is 2. Now, {observation} The identifier of the desklamp? Only Output a single number without any other words.’) found_desklamp = fdesklamp {answer}’ # Turn on the desklamp. observation = agent.turn_on(found_desklamp) # Check if the desklamp is turned on. assert turn on in observation, fError in [Step 6]: I cannot turn on {found_desklamp} in/on {receptacle}. {agent.report()}’ if start_from <= 7: print(”[Step 7] go to the receptacle where the book is and turn on the desklamp.”) observation = agent.goto(receptacle) # Check if the receptacle is closed. If so, open it. if closed in observation: observation = agent.open_receptacle(receptacle) # Turn on the desklamp. observation = agent.turn_on(found_desklamp) # Check if the desklamp is turned on. assert turn on in observation, fError in [Step 7]: I cannot turn on {found_desklamp} in/on {receptacle}. {agent.report()}’ if start_from <= 8: print(”[Step 8] examine the book with the desklamp.”) # Examine the book with the desklamp. observation = fYou examine the {found_book} with the {found_desklamp}.’ # Check if the book is examined with the desklamp. assert examine in observation, fError in [Step 8]: I cannot examine the {found_book} with the {found_desklamp}. {agent.report()}’ Success
Vanilla LLaMA3-8B-Instruct # Decomposition: General plan: I need to get a list of receptacles to find the book and take the book with me, then I get another list of receptacles to find the desklamp and examine the book with it. [Step 1] Get a list of receptacles where a book is likely to appear. [Step 2] Go to each receptacle in the list until seeing a book. [Step 3] Take the book from the receptacle. [Step 4] Get a list of receptacles where a desklamp is likely to appear. [Step 5] Go to each receptacle in the list until seeing a desklamp. [Step 6] Examine the book with the desklamp. # GPT Solution: def solution(agent, start_from=1): # General plan: I need to get a list of receptacles to find the book and take the book with me, then I get another list of receptacles to find the desklamp and examine the book with it. if start_from <= 1: print(”[Step 1] Get a list of receptacles where a book is likely to appear.”) # Ask the assistant to sort the receptacles based on the likelihood of finding a book. answer = ask(fGiven a list of receptacles, please sort them in descending order based on the likelihood of finding a book in each of them. The list of receptacles is: {agent.receptacles}. You should directly return a Python list.’) recep_to_check = literal_eval(answer) # Check if the returned list is not empty. assert recep_to_check, fError in [Step 1]: recep_to_check should not be empty. {agent.report()}’ if start_from <= 2: print(”[Step 2] Go to each receptacle in the list until seeing a book.”) for receptacle in recep_to_check: observation = agent.goto(receptacle) # Check if the receptacle is closed. If so, open it. if closed in observation: observation = agent.open_receptacle(receptacle) # Check if a book is in/on the receptacle. if book in observation: break # Check if a book is found in one of the receptacles. assert book in observation, fError in [Step 2]: There is no book in/on {recep_to_check}. {agent.report()}’ if start_from <= 3: print(”[Step 3] Take the book from the receptacle.”) # Get the identifier of the book from the observation. answer = ask(fFrom the observation, get the identifier of an object. For example, On the cabinet 1, you see a cloth 2, and a book 1. The identifier of cloth is 2. Now, {observation} The identifier of the book? Only Output a single number without any other words.’) found_book = fbook {answer}’ observation = agent.take(found_book, receptacle) # Check if the agent is holding the book. assert agent.holding == found_book, fError in [Step 3]: I cannot take {found_book} from the {receptacle}. {agent.report()}’ if start_from <= 4: print(”[Step 4] Get a list of receptacles where a desklamp is likely to appear.”) # Ask the assistant to sort the receptacles based on the likelihood of finding a desklamp. answer = ask(fGiven a list of receptacles, please sort them in descending order based on the likelihood of finding a desklamp in each of them. The list of receptacles is: {agent.receptacles}. You should directly return a Python list.’) recep_to_check = literal_eval(answer) # Check if the returned list is not empty. assert recep_to_check, fError in [Step 4]: recep_to_check should not be empty. {agent.report()}’ if start_from <= 5: print(”[Step 5] Go to each receptacle in the list until seeing a desklamp.”) for receptacle in recep_to_check: observation = agent.goto(receptacle) # Check if the receptacle is closed. If so, open it. if closed in observation: observation = agent.open_receptacle(receptacle) # Check if a desklamp is in/on the receptacle. if desklamp in observation: break # Check if a desklamp is found in one of the receptacles. assert desklamp in observation, fError in [Step 5]: There is no desklamp in/on {recep_to_check}. {agent.report()}’ if start_from <= 6: print(”[Step 6] Examine the book with the desklamp.”) # Examine the book with the desklamp. observation = agent.examine(found_book, found_desklamp # Check if the book is being examined with the desklamp. assert examine in observation, fError in [Step 6]: I cannot examine the book with the desklamp. {agent.report()}’ AttributeError: Agent object has no attribute examine’}, in step6: observation = agent.examine(found_book, found_desklamp)
Adaplanner def solution(agent, start_from=1): # General plan: I need to find the book and examine it with the desklamp. if start_from <= 1: print(”[Step 1] get a list of receptacles where a book is likely to appear.”) # Ask the assistant to sort the receptacles based on the likelihood of finding a book. answer = ask(fGiven a list of receptacles, please sort them in descending order based on the likelihood of finding a book in each of them. The list of receptacles is: {agent.receptacles}. You should directly return a Python list.’) recep_to_check = literal_eval(answer) # Check if the returned list is not empty. assert recep_to_check, fError in [Step 1]: recep_to_check should not be empty. {agent.report()}’ if start_from <= 2: print(”[Step 2] go to each receptacle in the list until finding a book”) for receptacle in recep_to_check: observation = agent.goto(receptacle) # Check if the receptacle is closed. If so, open it. if closed in observation: observation = agent.open_receptacle(receptacle) # Check if a book is in/on the receptacle. if book in observation: break # Check if a book is found in the receptacle. assert book in observation, fError in [Step 2]: There is no book in/on {recep_to_check}. {agent.report()}’ if start_from <= 3: print(”[Step 3] examine the book with the desklamp”) # Find the desklamp to use for examining the book. answer = ask(fGiven the observation, find the identifier of the desklamp. For example, On the desk 1, you see a lamp 1 and a book 2. The identifier of the lamp is 1. Now, {observation} What is the identifier of the desklamp? Only output a single number without any other words.’) found_desklamp = fdesklamp {answer}’ # Turn on the desklamp to examine the book. observation = agent.turn_on(found_desklamp) # Check if the desklamp is turned on. assert turn on in observation, fError in [Step 3]: Unable to turn on {found_desklamp} in/on {receptacle}. {agent.report()}’ AssertionError: Error in [Step 3]: Unable to turn on desklamp 0 in/on bed 1, in step3: assert turn on in observation, fError in [Step 3]: Unable to turn on {found_desklamp} in/on {receptacle}. {agent.report()}’
G.2 GSM-Hard
On the mathematical reasoning task, we compare M-Pilot with several strong baselines, including LLaMA3-8B-Instruct as the controller model, Chain-of-Thought (CoT) [51], and PaL [15]. In M-Pilot, the controller model is trained based on LLaMA3-8B-Instruct, and the black-box model employed is gpt-3.5-turbo-0125. Our results indicate that common issues with the baselines include incorrect code execution order and difficulties in translating questions into code. In contrast, M-Pilot significantly enhances the black-box model’s ability to solve mathematical problems through code generation.
Task Definition There are 5915126 students in Marissa’s class. Each student started the year with 10 pencils. After two months, 1/5 of the total pencils in class were used. At the end of the year, only 1/3 of the remaining pencils were left. How many pencils were left? True Answer: 15773669.333333334
M-Pilot # Decomposition: Lets break down this problem: 1. How many pencils were used in total? 2. How many pencils are left after using the total amount and then losing 2/3 of the remaining pencils? # GPT Solution: def solution(): ”””There are 5915126 students in Marissa’s class. Each student started the year with 10 pencils. After two months, 1/5 of the total pencils in class were used. At the end of the year, only 1/3 of the remaining pencils were left. How many pencils were left?””” # Initialization of variables num_students = 5915126 pencils_per_student = 10 pencils_used_fraction = 1/5 pencils_left_fraction = 1/3 # How many pencils were used in total? total_pencils_initial = num_students * pencils_per_student pencils_used = total_pencils_initial * pencils_used_fraction # How many pencils are left after using the total amount and then losing 2/3 of the remaining pencils? pencils_remaining = total_pencils_initial - pencils_used pencils_left = pencils_remaining * (1 - pencils_left_fraction) result = pencils_left return result Answer: 31547338.66666667, Wrong Answer # Second-Turn Decomposition: Lets break down this problem: 1. How many pencils were used in total? 2. How many pencils are left after using the total amount? Note: We dont need to consider losing 2/3 of the remaining pencils in the decomposition, as its a part of the calculation for the total pencils used. # Second-Turn GPT Solution: def solution(): ”””There are 5915126 students in Marissa’s class. Each student started the year with 10 pencils. After two months, 1/5 of the total pencils in class were used. At the end of the year, only 1/3 of the remaining pencils were left. How many pencils were left?””” # Initialization of variables num_students = 5915126 pencils_per_student = 10 pencils_used_fraction = 1/5 pencils_left_fraction = 1/3 # How many pencils were used in total? total_pencils_initial = num_students * pencils_per_student pencils_used = total_pencils_initial * pencils_used_fraction # How many pencils are left after using the total amount and then losing 2/3 of the remaining pencils? pencils_remaining = total_pencils_initial - pencils_used pencils_left = pencils_remaining * (1 - pencils_left_fraction) result = pencils_left return result Answer: 15773669.333333332, Correct Answer
Vanilla LLaMA3-8B-Instruct # Decomposition: Lets break down this problem: 1. How many pencils were used in total? 2. How many pencils were left after the first two months? 3. How many pencils were left at the end of the year? # GPT Solution: def solution(): ”””There are 5915126 students in Marissa’s class. Each student started the year with 10 pencils. After two months, 1/5 of the total pencils in class were used. At the end of the year, only 1/3 of the remaining pencils were left. How many pencils were left?””” # Initialization of variables num_students = 5915126 pencils_per_student = 10 pencils_used_fraction = 2/5 pencils_left_fraction = 1/3 # How many pencils were used in total? total_pencils_initial = num_students * pencils_per_student pencils_used_total = total_pencils_initial * pencils_used_fraction # How many pencils were left after the first two months? pencils_remaining_after_two_months = total_pencils_initial - pencils_used_total # How many pencils were left at the end of the year? pencils_left = pencils_remaining_after_two_months * pencils_left_fraction result = pencils_left return result Answer: 11830252.0, Wrong Answer
PaL def solution(): ”””There are 5915126 students in Marissa’s class. Each student started the year with 10 pencils. After two months, 1/5 of the total pencils in class were used. At the end of the year, only 1/3 of the remaining pencils were left. How many pencils were left?””” num_students = 5915126 pencils_per_student = 10 total_pencils_initial = num_students * pencils_per_student pencils_used = total_pencils_initial // 5 pencils_remaining = total_pencils_initial - pencils_used pencils_left_end_of_year = pencils_remaining // 3 result = pencils_left_end_of_year return result Answer: 15773669.0, Wrong Answer
CoT First, we need to find out how many pencils were used after two months. Since 1/5 of the total pencils were used, the number of pencils used is 1/5 * (5915126 * 10) = 1183025.2 pencils. Next, we need to find out how many pencils were left at the end of the year. Since only 1/3 of the remaining pencils were left, the number of pencils left is 2/3 * (5915126 * 10 - 1183025.2) = 3156750.8 pencils. Since we cant have a fraction of a pencil, we need to round down to the nearest whole number. Therefore, the number of pencils left is 3156750 pencils. So the answer is 3156750. Answer: 3156750, Wrong Answer
G.3 LaMP
We utilize case study tables to illustrate the performance improvements achieved through the intermediate generations of the controller model. Specifically, we present detailed case studies for LaMP-2M, LaMP-2N, and LaMP-3 in Table 20. In these analyses, the controller model is Llama-3-8B-Instruct, and the environment model is gpt-4o-mini. As demonstrated in Table 20, M-Pilot enhances the summarization capabilities of the controller model. Common errors in PAG, such as inaccurate summaries or improper utilization of retrieved user history, are mitigated by M-Pilot, which enables the controller model to produce more accurate generations through interaction and optimization.
Table 20: Case Study for LaMP. In the ”Target” column, we present the ground-truth categorization. The ”Gen” column displays the final answer generated by the black-box model corresponding to each intermediate generation. The ”Score” column indicates whether the generated answer (”Gen”) matches the target categorization (”Target”). In the ”Method” column, ”Ours” refers to M-Pilot, while ”PAG” stands for Profile Augmented Generation.
Task Input Question Intermediate Generation Method Target Gen Score 2M Which tag does this movie relate to among the following tags? A ticking-time-bomb insomniac and a slippery soap salesman channel… most popular tag: ”dystopia”, ”fantasy”, ”comedy”, ”violence” Ours violence violence 1 Here are the most popular tags for the user: dystopia, fantasy, comedy PAG violence dystopia 0 2N Which category does this article relate to among the following categories? The suspect, Akayed Ullah, was the most seriously hurt in the rush-hour blast… most popular category: politics, crime, entertainment, women, business, sports. Ours crime crime 1 Based on the articles you provided, the most popular category written by this journalist is: politics. PAG crime politics 0 3 What is the score of the following review on a scale of 1 to 5? After almost 20 years in and around MIT, I’ve encountered only two great MIT books: (1) A.R. Gurney’s out-of-print novel The Snow Ball (correction: it is Entertaining Strangers); (2) Pepper White’s book… Based on this user’s past reviews, the most common positive score is: 4, with 4 reviews out of 8 receiving a score of 4. The most common negative score is: 1, with 4 reviews out of 8 receiving a score of 1. Ours 4 4 1 Based on the reviews, the most common positive score is 5, and the most common negative score is 1. PAG 4 5 0
Appendix H Prompt Templates
H.1 ALFWorld
@@
Following Adaplanner [41], we implement a code-style prompt for M-Pilot, which can be divided into the following sections:
@@
The <high_level_planning> prompt is used to instruct M-Pilot to break the current task down into multiple subtasks, where <decompose> is replaced by a standard task decomposition process, and <receptacle_list> is substituted by the list of interactive receptacles provided by the task environment. Finally, <task> is replaced by the task description, expressed in natural language.
<high_level_planning> Prompt # Decompose the task into steps. First give a general plan of how you would solve the task, then for each step you plan to take, mark with ’[Step $x$]’. # Here is an example of a decomposition to the task: # define environment receptacles = [’diningtable 1’,’drawer 2’, drawer 1’, sinkbasin 1’, toilet 1’, sidetable 2’, sidetable 1’, cabinet 1’, countertop 1’, microwave 1’, fridge 1’] <decompose> # Here is the actual task. # define environment receptacles = <receptacle_list> # <task> # here is a decomposition:
Multi-Turn Planning.
The <multi_turn_planning> prompt is used to guide M-Pilot to reflect on errors in the previous decomposition and re-break the current task into multiple subtasks. Here, <predecompose> is replaced with the high-level plan from the previous turn, while all other elements retain the same meaning as before.
<multi_turn_planning> Prompt # Decompose the task into steps. First give a general plan of how you would solve the task, then for each step you plan to take, mark with ’[Step $x$]’. # Here is a successful example of a decomposition to the task: # define environment receptacles = [’diningtable 1’,’drawer 2’, drawer 1’, sinkbasin 1’, toilet 1’, sidetable 2’, sidetable 1’, cabinet 1’, countertop 1’, microwave 1’, fridge 1’] <decompose> # Here is the actual task. # define environment receptacles = <receptacle_list> # <task> # Here are the decomposition steps you previously generated for the task. <predecompose> # However, you made a mistake in the decomposition above because of lack of understanding of the task. Referring to the successful example, please correct the error, if any, and rewrite the decomposition. # here is a decomposition:
@@
The <low_level_execution> prompt is used to instruct the black box model to generate a specific solution based on the problem and the plan provided by M-Pilot. <basic_info> defines the agent and admissible actions on Alfworld and can be found in [41]. The <example> is replaced with a combination of planning and expert trajectory, while the meanings of <receptacle_list> and <task> remain consistent with the previous description. <decomposition> represents the high-level plan provided by M-Pilot for the current task.
<low_level_execution> Prompt <basic_info> # Now complete the function solution() below to solve the task by composing the agents methods to interact with the environment. # First give a general plan of how you would solve the task, mark with # General Plan’. Then for each step you plan to take, 1) mark with ’[Step $x$]’, 2) give a reason why you think it is a good step to take 3) write an assertion to check if the step is successful. # Here is an example of a solution to the task: # define environment and agent receptacles = [’diningtable 1’,’drawer 2’, drawer 1’, sinkbasin 1’, toilet 1’, sidetable 2’, sidetable 1’, cabinet 1’, countertop 1’, microwave 1’, fridge 1’] agent = Agent(receptacles) <example> # Here is the actual task. # define environment and agent receptacles = <receptacle_list> agent = Agent(receptacles) # <task> # here is a decomposition: <decomposition> # here is a solution:
@@
In ALFWorld, there are six types of tasks: Pick, Clean, Heat, Cool, Examine, and Pick two. For each type, we collect a reasonable high-level planning approach, allowing M-Pilot to reference them. These six planning samples are presented as follows:
@@
<planning_sample_pick> Prompt # Your task is to: put soapbar on countertop. # here is a decomposition: # General Plan: I need to get a list of receptacles where the soapbar is likely to appear, and then go to each receptacle in the list until seeing a soapbar. Then I can put get the identifier of the soapbar and take it. Finally I can go to the countertop and put the soapbar. # [Step 1] get a list of receptacles where the soapbar is likely to appear. # [Step 2] go to each receptacle in the list until seeing a soapbar. # [Step 3] identify the soapbar I juts found and take it. # [Step 4] go to a countertop and put the soapbar on it.
@@
<planning_sample_clean> Prompt # Your task is to: put a clean lettuce in diningtable / clean a lettuce and put it in diningtable. # here is a decomposition: # General plan: I need to get a list of receptacles to find the lettuce, take the lettuce to the sinkbasin, clean it and put it in a diningtable. # [Step 1] get a list of receptacles where the lettuce is likely to appear. # [Step 2] go to each receptacle in the list until seeing a lettuce. # [Step 3] identify the lettuce I just found and take it. # [Step 4] go to a sinkbasin to clean the lettuce. # [Step 5] go to a diningtable and put the lettuce on it.
@@
<planning_sample_heat> Prompt # Your task is to: put a hot lettuce in diningtable / heat some lettuce and put it in diningtable. # here is a decomposition: # General plan: I need to get a list of receptacles to find the lettuce, take the lettuce to the microwave, heat it and put it in a diningtable. # [Step 1] get a list of receptacles where the lettuce is likely to appear. # [Step 2] go to each receptacle in the list until seeing a lettuce. # [Step 3] identify the lettuce I juts found and take it. # [Step 4] go to a microwave to heat the lettuce. # [Step 5] go to a diningtable and put the lettuce on it.
@@
<planning_sample_cool> Prompt # Your task is to: put a cold lettuce in diningtable / cool some lettuce and put it in diningtable. # here is a decomposition: # General plan: I need to get a list of receptacles to find the lettuce, take the lettuce to the fridge, cool it and put it in a diningtable. # [Step 1] get a list of receptacles where the lettuce is likely to appear. # [Step 2] go to each receptacle in the list until seeing a lettuce. # [Step 3] identify the lettuce I juts found and take it. # [Step 4] go to a fridge to cool the lettuce. # [Step 5] go to a diningtable and put the lettuce on it.
@@
<planning_sample_examine> Prompt # Your task is to: look at the bowl under the desklamp / examine the bowl with the desklamp # here is a decomposition: # General plan: I need to get a list of receptacles to find the bowl and take the bowl with me, then I get another list of receptacles to find the desklamp and turn it on. # [Step 1] get a list of receptacles where a bowl is likely to appear. # [Step 2] go to each receptacle in the list until seeing a pen. # [Step 3] take the bowl from the receptacle. # [Step 4] get a list of receptacles where a desklamp is likely to appear. # [Step 5] go to each receptacle in the list until seeing a desklamp. # [Step 6] turn on desklamp.
@@
<planning_sample_picktwo> Prompt # Your task is to: put two cellphone in cabinet / find two cellphone and put them in cabinet # here is a decomposition: # General plan: I need to get a list of receptacles to find the two cellphones, find and take the first cellphone and put it in a cabinet, then find and take the second cellphone and put it in the cabinet. # [Step 1] get a list of receptacles where a cellphone is likely to appear. # [Step 2] go to each receptacle in the list until seeing a cellphone. # [Step 3] identify the first cellphone found and take it. # [Step 4] go to a cabinet and put the first cellphone found on it. # [Step 5] go to each of the remaining receptacle in the list until seeing a second cellphone. # [Step 6] identify the second cellphone I just found and take it. # [Step 7] go to a cabinet and put the second cellphone found on it.
@@
Our execution sample is based on the prompt structure from [41], with the key distinction being the incorporation of the planning component. In this setup, <decompose> is substituted with the task-specific planning sample, <execution> is replaced by the expert samples from [41], and the definition of <task> remains unchanged from the previous description.
<execution_sample_template> Prompt # <task> # <decompose> # <execution>
@@
To implement close-loop refinement during the inference stage, we follow the approach from [41] and introduce several prompts: a <code_check> prompt to identify and fix any syntax errors during execution generation, a <refinement> prompt to address refinement in case of assertion errors, and a <start_from> prompt to determine the starting point for the new solution after revising the plan. Detailed descriptions of these prompts can be found in [41].
H.2 GSM-Hard
Following PAL framework [15], we implement a code-based framework to solve mathematical problems on GSM-Hard, which is primarily divided into two steps: M-Pilot breaks down the mathematical problem into sub-problems, and the black-box model converts each sub-problem into a code block.